You find an unmarked blue laser on your way to physics class. When you get to class you realize that you can determine the wavelength of the laser by doing a double-slit experiment. Shining the laser through a double slit with a separation of 0.351 mm and projecting the interference pattern on the wall 2.17 m away, the first bright fringe is 2.85 mm from the center of the pattern. What is the wavelength?
nm
Answer #1

This problem is based on double slit experiment,

As we know, Fringe width

\beta=\dfrac{D \lambda }{d} Here, D: Distance between the screen and slit

d: slit width

  \lambda: Wavelength of light used

Also we know first maxima will be at a distance of 1\beta from the center.

therefore using the above condition, we get-

2.85*10^{-3}=1\beta

Using expression for \beta we get,

2.85*10^{-3}=1\dfrac{D \lambda}{d}

\lambda=2.85*10^{-3}\dfrac{d}{D*1}

after using given values of D and d,

\lambda=2.85*10^{-3}\dfrac{0.351*10^{-3}}{2.17*1}

\lambda=460.99*10^{-9}m

\lambda=461 nm

Thankyou.